# Learning Object Compliance via Young’s Modulus from Single Grasps using Camera-Based Tactile Sensors

### v3

@@ Abstract
Compliance is a useful parametrization of tactile information that humans often utilize in manipulation tasks. It can be used to inform low-level contact-rich actions or characterize objects at a high-level. In robotic manipulation, existing approaches to estimate compliance have struggled to generalize across both object shape and material. Using camera-based tactile sensors, proprioception, and force measurements, we present a novel approach to estimate object compliance as Young’s modulus $E$ from parallel grasps. We evaluate our method over a novel dataset of 285 common objects, including a wide array of shapes and materials with Young’s moduli ranging from 5.0 kPa to 250 GPa. Combining analytical and data-driven approaches, we develop a hybrid system using a multi-tower neural network to analyze a sequence of tactile images from grasping. This system is shown to estimate the Young’s modulus of unseen objects within an order of magnitude at 74.2% accuracy across our dataset. This is an improvement over purely analytical and data-driven baselines which exhibit 28.9% and 65.0% accuracy respectively. Importantly, this estimation system performs irrespective of object geometry and demonstrates increased robustness across material types.
@@ I Introduction
To address these limitations and move towards more generalized compliance estimation, we propose a hybrid approach that combines analytical modeling with learning-based methods to estimate an object’s compliance as Young’s modulus $E$ from a single parallel grasp. By incorporating analytical models, we improve generalization across materials, while using a learned CNN to account for geometric complexities. We choose to label compliance as Young’s modulus to align with analytical models and so that our system produces estimations on a universal scale.
@@ III-B Analytical Contact Models
With camera-based tactile sensors, we can increase the fidelity of contact models by measuring contact area $A$ and depth $d$ [8]. Contact area $A$ is determined by masking tactile depth images with a constant threshold of 0.1mm. This threshold was chosen empirically to cancel out potential noise. Without tactile sensing, we could not measure depth ($d(t)=0$) and would need to assume a constant contact area ($A(t)=A_{\text{sensor}}$). We calculate contact stress $\sigma$ and object strain $\epsilon$, following Eq. 1 and Eq. 2. This model assumes symmetry between parallel fingers. We define $t=0$ to be the moment of first contact, where $F(t=0)$ must be greater than 0.75 N.
@@
Hertzian Contact: Hertzian contact theory models the deformation of elastic bodies in contact with each other. In a simple case, contact is assumed to be without friction or adhesion [10, 9]. Unlike previous methods, Hertzian mechanics closely analyzes geometry through contact. It has poreviously been applied to model soft contact in robotics with camera-based tactile sensors [23, 25, 22, 13].
@@ VI Results
Furthermore, we utilize this reduced-domain training configuration to evaluate our architecture against previously-developed Shore hardness estimation models. We can only compare between predictions in this domain because previously-developed models have only been trained for use with rubber objects. Converting our model’s predictions from Young’s modulus to Shore 00 hardness [43], we find that our model reduces root mean square error (RMSE) compared to Yuan et al. [12] from 10.3 to 5.9 over simple-shaped objects and from 18.2 to 14.3 over arbitrarily-shaped objects, in units of Shore hardness. This observed improvement is likely due to our method’s integration of analytical estimates alongside tactile image-based CNN models, whereas previously methods were fully data-driven.

### v4

@@ Abstract
Compliance is a useful parametrization of tactile information that humans often utilize in manipulation tasks. It can be used to inform low-level contact-rich actions or characterize objects at a high-level. In robotic manipulation, existing approaches to estimate compliance have struggled to generalize across both object shape and material. Using camera-based tactile sensors, proprioception, and force measurements, we present a novel approach to estimate object compliance as Young’s modulus $E$ from parallel grasps. We evaluate our method over a novel dataset of 285 common objects, including a wide array of shapes and materials with Young’s moduli ranging from 5.0 kPa to 250 GPa. Combining analytical and data-driven approaches, we develop a hybrid system using a multi-tower neural network to analyze a sequence of tactile images from grasping. This system is shown to estimate the Young’s modulus of unseen objects within an order of magnitude at 74.2% accuracy across our dataset. This is an improvement over purely analytical and data-driven baselines which exhibit 28.9% and 65.0% accuracy respectively. Importantly, this estimation system performs irrespective of object geometry and demonstrates increased robustness across material types. Code is available on GitHub and collected data is available on HuggingFace.
@@ I Introduction
To address these limitations and move towards more generalized compliance estimation, we propose a hybrid approach that combines analytical modeling with learning-based methods to estimate an object’s compliance as Young’s modulus $E$ from a single parallel grasp. By incorporating analytical models, we improve generalization across materials, while using a learned CNN to account for geometric complexities. We label compliance as Young’s modulus to align with analytical models and produce estimations on a universal scale.
@@ III-B Analytical Contact Models
With camera-based tactile sensors, we can increase the fidelity of contact models by measuring contact area $A$ and depth $d$ [8]. Contact area $A$ is determined by masking tactile depth images with a constant threshold of 0.1mm. This threshold was chosen empirically to cancel out potential noise. Without tactile sensing, we could not measure depth ($d(t)=0$) and would need to assume a constant contact area ($A(t)=A_{\text{sensor}}$). We calculate contact stress $\sigma$ and object strain $\epsilon$, following Eq. 1 and Eq. 2. This model assumes symmetry between parallel fingers. We define $t=0$ to be the moment of first contact, where $F(t=0)$ must be greater than an impirically chosen threshold of 0.75 N.
@@
Hertzian Contact: Hertzian contact theory models the deformation of elastic bodies in contact with each other. In a simple case, contact is assumed to be without friction or adhesion [10, 9]. Unlike previous methods, Hertzian mechanics closely analyzes geometry through contact. It has been applied to model soft contact in robotics with camera-based tactile sensors [23, 25, 22, 13].
@@ VI Results
Furthermore, we utilize this reduced-domain training configuration to evaluate our architecture against previously-developed Shore hardness estimation models. We can only compare between predictions in this domain because previously-developed models have only been trained for use with rubber objects. Converting our model’s predictions from Young’s modulus to Shore 00 hardness [43], we find that our model reduces root mean square error (RMSE) compared to Yuan et al. [12] from 10.3 to 5.9 over simple-shaped objects and from 18.2 to 14.3 over arbitrarily-shaped objects, in units of Shore hardness. This observed improvement is likely due to our method’s integration of analytical estimates alongside tactile image-based CNN models, whereas previous methods were fully data-driven.